# Differentially Private Representation Learning via Image Captioning

### v1

@@ Abstract
Differentially private (DP) machine learning is considered the gold-standard solution for training a model from sensitive data while still preserving privacy. However, a major barrier to achieving this ideal is its sub-optimal privacy-accuracy trade-off, which is particularly visible in DP representation learning. Specifically, it has been shown that under modest privacy budgets, most models learn representations that are not significantly better than hand-crafted features. In this work, we show that effective DP representation learning can be done via image captioning and scaling up to internet-scale multimodal datasets. Through a series of engineering tricks, we successfully train a DP image captioner (DP-Cap) on a 233M subset of LAION-2B from scratch using a reasonable amount of computation, and obtaining unprecedented high-quality image features that can be used in a variety of downstream vision and vision-language tasks. For example, under a privacy budget of $\varepsilon=8$, a linear classifier trained on top of learned DP-Cap features attains $65.8\%$ accuracy on ImageNet-1K, considerably improving the previous SOTA of $56.5\%$. Our work challenges the prevailing sentiment that high-utility DP representation learning cannot be achieved by training from scratch.
[…]
Figure 1: (a) Few-shot ImageNet-1K linear probe accuracy comparison between DP-Cap (ours) and ViP (Yu et al., 2023) (previous SOTA). DP-Cap learns better image representations using the same training data and privacy budget, and considerably surpasses synthetic initialization (syn). (b) Compositional understanding evaluation on the ARO benchmark (Yuksekgonul et al., 2022). DP-Cap performance is close to non-private Cap and outperforms non-private CLIP. (c) Captions generated by DP-Cap on images from the MS-COCO 2017 (Lin et al., 2015) test set.
@@ 1 Introduction
Differentially private (DP; Dwork et al. (2006)) model training is an effective strategy for privacy-preserving ML on sensitive data. For most optimization-based learning algorithms, DP-SGD (Song et al., 2013; Abadi et al., 2016) can be readily applied to obtain models with rigorous DP guarantee. Regrettably, DP training has also been marred by a sub-optimal privacy-utility trade-off, with model utility severely lagging behind their non-private counterpart (Jayaraman and Evans, 2019; Tramer and Boneh, 2020; Kurakin et al., 2022). At the core of this unfavorable trade-off is the difficulty of DP representation learning. Tramer and Boneh (2020) showed that when DP training from scratch under a low-to-moderate privacy budget, most models learn representations with a quality worse than even handcrafted features. These observations naturally lead to the research question: “How does one learn useful representations with DP training?”
One plausible reason for the failure of prior attempts at DP representation learning is the lack of training data. Indeed, DP limits the information content of each training sample via the privacy budget $\varepsilon$, inducing a privacy-accuracy-sample size tradeoff; thus a substantially larger training dataset is required to extract the same amount of information to train the model. As the vast majority of prior work only utilize small to moderate scale classification datasets such as CIFAR-10 (Krizhevsky et al., 2009) and ImageNet (Deng et al., 2009b), the amount of training data is simply insufficient for learning high-quality representations under DP (Tramer and Boneh, 2020). Yu et al. (2023) made partial progress towards this through self-supervised learning (SSL) on internet-scale data. By training a masked autoencoder (MAE; He et al. (2022)) using DP-SGD on a 233M subset of the LAION-2B dataset (Schuhmann et al., 2022), the model learned image representations that are on-par with non-private AlexNet (Krizhevsky et al., 2012) trained on ImageNet—the first deep learning model to outperform handcrafted features and a major cornerstone for representation learning. However, the MAE objective promotes the model to learn extraneous details in the image that may not be helpful for obtaining generalizable representations, limiting the potential of this approach for DP.
@@
The image representations learned by DP-Cap also exhibit strong performance for multimodal tasks that require alignment of image and text features, the first occurrence for models trained from scratch with DP; see Figure 1(b). As a qualitative evaluation, we also use the trained DP-Cap model to caption several images from the MS-COCO 2017 (Lin et al., 2015) test set in Figure 1(c) and Appendix 7.3. The resulting captions are grammatically correct and semantically coherent, while (close to) accurately describing contents of the image; this is interesting because our model has only been exposed to language supervision from LAION, which are far from being flawless. Our results suggest that DP training on internet-scale multimodal datasets can be a viable approach for obtaining high-utility learned representations.
@@ 2 Background and Related Work
Vision-language pre-training. Many modern ML datasets such as Conceptual Captions (Changpinyo et al., 2021), LAION (Schuhmann et al., 2021) and DataComp (Gadre et al., 2023) consist of aligned image-text pairs where the image and text contain roughly similar semantic information. One can leverage the aligned nature of the training data to pre-train vision-language models (VLMs) that connect the two modalities, whose representations perform more general multi-modal tasks. Contrastive learning-based techniques such as CLIP (Radford et al., 2021) and BLIP (Li et al., 2022) are also applicable for pre-training VLMs. Doing so not only learns high-quality image and text representations but also introduces new multi-modal capabilities such as cross-modal retrieval and zero-shot prediction (Radford et al., 2021). Recent work by Tschannen et al. (2023) shows an image captioning approach (predicting text captions from images) is a viable alternative to contrastive learning and can lead to models with robust performance.
@@
$\widetilde{\mathbf{g}}_{k}:=\frac{1}{B}\left[\sum_{i\in\mathcal{B}_{k}}\text{clip}_{C}\left(\nabla_{\bm{\theta}}\ell_{i}(\bm{\theta}_{k})\right)+\mathcal{N}\left(0,C^{2}\sigma^{2}\mathbf{I}\right)\right],$ (2)
where $\ell_{i}(\bm{\theta})$ is the per-sample loss function evaluated at sample $\mathbf{x}_{i}$. We also use the term “DP-SGD” loosely to refer to the category of gradient-based optimization algorithms that operate on the noisy gradient, e.g., Adam (Kingma and Ba, 2014). The privacy analysis of DP-SGD relies on composition of multiple steps. One particularly powerful analysis framework amenable to such compositions relies on a variant of DP called Rényi differential privacy (RDP) (Mironov, 2017). An advantage of RDP is its additive composition property, where the privacy guarantees of a sequence of mechanisms can be combined with amplification by subsampling (Wang et al., 2019) and then translated to $(\varepsilon,\delta)$-DP (Balle et al., 2020; Gopi et al., 2021). In this work, we adopt this accounting technique.
Scaling up DP-SGD training. DP training is a theoretically and empirically proven remedy against unintended training data memorization. Even models with large $\varepsilon$ (e.g., $\varepsilon=100$) can empirically defend against privacy attacks (Carlini et al., 2021; Guo et al., 2023). Despite its great appeal, DP training also carries a significant drawback of large drop in model utility (Abadi et al., 2016; Tramer and Boneh, 2020). For example, the SOTA performance on ImageNet when training from scratch with a DP guarantee of $\varepsilon=8$ is 39.2% (Sander et al., 2023); in comparison, the non-private performance on ImageNet when training from scratch can reach 88% (Touvron et al., 2022) or higher. This degradation in model utility also translates to poorly learned representations, as Tramer and Boneh (2020) showed that even handcrafted features can rival ones learned through DP training.
@@ 3.1 DP Representation Learning via Image Captioning
Why is vision-language pre-training suitable? Given image-text aligned datasets, prior works (Radford et al., 2021; Li et al., 2022; Tschannen et al., 2023) showed that pre-training using language supervision is an appealing option for non-private representation learning. We hypothesize that this is true for DP representation learning as well. Compared to image-only supervision, language supervision contains a more condensed summary of the image content, allowing the model to ignore irrelevant details such as background and focus on objects of interest and their relationships. This is especially helpful for DP since the model needs to extract as much useful information as possible from each sample given the privacy budget $\varepsilon$. Captioning could thus enhance the privacy-utility-sample size trade-off in DP, considering it requires less information per sample.
In addition, we show that vision-language pre-training supports a very large batch size, much larger than what is typically used in image-only pre-training (Radford et al., 2021; Li et al., 2022; Yu et al., 2022). This subtle aspect is in fact crucial for reducing the effective noise in DP-SGD (Li et al., 2021), which allows the model parameters to converge to a stable solution with lower training loss (see Section 3.2).
@@
Unlike contrastive learning, the image captioning approach (Sariyildiz et al., 2020; Desai and Johnson, 2021; Tschannen et al., 2023) aligns well with DP-SGD training. Specifically, an image captioner is trained to predict captions based on their corresponding images. The training objective of the image captioner for one image-text pair $[\mathbf{x}^{\text{img}},\mathbf{z}^{\text{text}}]$ is to minimize over $\bm{\theta}:=\{\bm{\theta}_{\textsf{enc}},\bm{\theta}_{\textsf{dec}}\}$ the following loss:
$L_{\text{Cap}}([\mathbf{x}^{\text{img}},\mathbf{z}^{\text{text}}];\bm{\theta}):=\frac{1}{T}\sum_{t=0}^{T-1}\ell_{\text{CE}}\Big{(}{z}^{\text{text}}_{t+1},\varphi\big{(}\underbrace{\psi(\mathbf{x}^{\text{img}};\bm{\theta}_{\textsf{enc}})}_{\text{image embedding}},\underbrace{{z}^{\text{text}}_{1},\ldots,{z}^{\text{text}}_{t}}_{\text{first $t$ tokens}};\bm{\theta}_{\textsf{dec}}\big{)}\Big{)},$ (3)
where $\mathbf{z}^{\text{text}}$ denotes the caption token sequence $\{{z}^{\text{text}}_{1},\ldots,{z}^{\text{text}}_{T}\}$ and the image captioner consists of two parts: the image encoder $\psi(\cdot;\bm{\theta}_{\textsf{enc}})$ and the text decoder $\varphi(\cdot;\bm{\theta}_{\textsf{dec}})$. The rationale behind the design of equation 3 is that the image encoder maps the input image $\mathbf{x}^{\text{img}}$ to an embedding vector, and the text decoder takes the image embedding $\psi(\mathbf{x}^{\text{img}};\bm{\theta}_{\textsf{enc}})$ and the first $t$ caption tokens $\{{z}^{\text{text}}_{1},\ldots,{z}^{\text{text}}_{t}\}$ as inputs and predicts the next caption token ${z}_{t+1}^{\text{text}}$. Both the encoder and decoder are trained to maximize the log-likelihood of the correct next token. Equation 3 corresponds to the loss function for an image-text pair $[\mathbf{x}^{\text{img}},\mathbf{z}^{\text{text}}]$; summing over all the samples in a batch gives the complete empirical loss in an additive form, which is directly compatible with DP-SGD. We discuss the sense of the privacy guarantees in Appendix 6.1.
@@ 3.2 Strategy for Effective DP Training
Figure 2: Impact of synthetic initialization on the DP-Cap model. The learned image representation benefits substantially from initializing on the Shaders21k dataset. The accuracy gap between DP-Cap (random init) and DP-Cap (syn init) can be as large as 24% for ImageNet linear probing.
@@
(b) Performance improves consistently with the batch size
@@
Figure 3: (a) We fix the effective noise $\sigma/B=5.6\times 10^{-7}$ (corresponding to our (B, $\sigma$) = (1.3M, 0.728)) and show that the loss is remarkably consistent across different batch sizes, allowing us to effectively scale up batch size to improve the SNR. (b) Performance from 4 sets of parameters that provide $\varepsilon=8$, with constant number of steps 5708. From batch size 98k (used in ViP (Yu et al., 2023)), to our 1.3M batch size. In contrast to ViP, DP-Cap successfully leverages the better SNR and learns features that achieve substantially better 10-shot accuracy on ImageNet even compared to a non-private MAE (He et al., 2022) trained on the same dataset (see Appendix 6.1).
@@
Sufficient number of training steps. We address this challenge via synthetic pre-training. We first observe that image representations learned by DP-Cap (random init) outperform those of ViP (random init) as evidenced in table 9 in App. 7. Interestingly, Yu et al. (2023) have shown that synthetic images consisting of only textures can provide a better initialization for their reconstruction-based model without any privacy risk. With this initialization, the model can focus on learning dataset-specific properties rather than low-level image properties such as edge detectors, therefore expending privacy budget in a more optimal manner. We adapt this technique of pre-training on the Shaders21k dataset (Baradad et al., 2022) for initialization (see Appendix 6.1 for details) and observe that it is even more effective with DP-Cap compared to ViP. As shown in Fig. 2, our DP-Cap (syn init) improves over DP-Cap (random init) by more than 24% on ImageNet-1k linear probing. It also improves over synthetic initialization alone (Syn-init) by more than 14% on ImageNet-1k linear probing, whereas the gain in ViP is smaller than 6%.
@@
Using extreme batch sizes to reduce effective noise. Li et al. (2021) first showed that increasing the batch size in DP-SGD often improves the privacy-utility trade-off. This is because the effective noise added to the average gradient has magnitude $\sigma/B$ (cf. equation 2), and that increasing $B$, rather than decreasing $\sigma$, results in better privacy guarantees according to conventional accounting techniques (Bun and Steinke, 2016; Dwork and Rothblum, 2016; Mironov et al., 2019; Sander et al., 2023). However, under supervised learning, increasing the batch size beyond a certain limit can lead to training instability under both private and non-private training. Specifically, Sander et al. (2023, 2024) observed that when training a classifier from scratch with DP-SGD on ImageNet, at a fixed number of steps $S$ and fixed effective noise $\sigma/B$, the performance decreases significantly when the batch size becomes too large: for $B\in[128,16384]$, a drop of 10% in top-1 accuracy was observed.
@@
Intriguingly, we find that vision-language pre-training on internet-scale datasets can tolerate extreme batch sizes, e.g. $B=1$M. In Figure 3(a), we compare the loss behaviors when scaling the batch size for DP-Cap. We fix the effective noise $\sigma/B$ while varying the batch size. In stark contrast to the previous observation from Sander et al. (2023), the loss trajectory is identical across different batch sizes. With this observation, we are able to successfully scale up the batch size for DP-Cap to as large as $B=1.3$M, achieving an effective noise of $5.6\times 10^{-7}$, almost 10 times smaller than the effective noise of ViP in Yu et al. (2023). Training DP-Cap under such a small effective noise allows it to extract information from the training dataset more efficiently under the DP constraint. In Figure 3, we show that with $B=1.3$M, the representation learned by DP-Cap even outperforms non-private MAE trained on the same dataset.
@@
Figure 4: Number of GPU hours to train DP-Cap for one epoch on 233M samples. For the Large model, we achieve a close to 5×5$\times$5 × reduction.
@@
Method Model Tiny Small Base Large functorch 36 20 12 5 Ghost norm 350 230 180 64 Ghost norm+AMP 420 310 225 90 Table 1: Max physical batch size under different per-sample gradient computation methods.
Figure 4: Number of GPU hours to train DP-Cap for one epoch on 233M samples. For the Large model, we achieve a close to 5×5$\times$5 × reduction.
@@
Improving training pipeline efficiency. DP training is known to be computationally inefficient due to factors such as per-sample gradient computation (Lee and Kifer, 2020; Li et al., 2021). Training on internet-scale datasets using extreme batch sizes further complicates this issue. For example, a naive implementation of DP-Cap with per-sample gradient computation using functorch would take approximately 61 days(!) on 128 NVIDIA V100 GPUs. We made significant efficiency improvements to the training pipeline using two techniques: ghost norm (Li et al., 2021) and the automatic mixed precision (AMP) package in PyTorch. Combining these two techniques with DP-SGD requires careful considerations to ensure both a correct DP guarantee as well as numerical stability. We detail the implementation of these two techniques in Appendix 6.2.

### v2

@@ Abstract
Differentially private (DP) machine learning is considered the gold-standard solution for training a model from sensitive data while still preserving privacy. However, a major barrier to achieving this ideal is its sub-optimal privacy-accuracy trade-off, which is particularly visible in DP representation learning. Specifically, it has been shown that under modest privacy budgets, most models learn representations that are not significantly better than hand-crafted features. In this work, we show that effective DP representation learning can be done via image captioning and scaling up to internet-scale multimodal datasets. Through a series of engineering tricks, we successfully train a DP image captioner (DP-Cap) on a 233M subset of LAION-2B from scratch using a reasonable amount of computation, and obtaining unprecedented high-quality image features that can be used in a variety of downstream vision and vision-language tasks. For example, under a privacy budget of $\varepsilon=8$ for the LAION dataset, a linear classifier trained on top of learned DP-Cap features attains $65.8\%$ accuracy on ImageNet-1K, considerably improving the previous SOTA of $56.5\%$. Our work challenges the prevailing sentiment that high-utility DP representation learning cannot be achieved by training from scratch. Code is available at https://github.com/facebookresearch/dpcap.
Machine Learning, ICML
Figure 1: (a) Few-shot ImageNet-1K linear probe accuracy comparison between DP-Cap (ours) and ViP (Yu et al., 2023) (previous SOTA). DP-Cap learns better image representations using the same training data and privacy budget, and considerably surpasses synthetic initialization (syn). The privacy budget $\varepsilon$ is for the LAION dataset, and the linear classifiers are trained without DP. (b) Compositional understanding evaluation on the ARO benchmark (Yuksekgonul et al., 2022). DP-Cap performance is close to non-private Cap and outperforms non-private CLIP. (c) Captions generated by DP-Cap on images from the MS-COCO 2017 (Lin et al., 2015) test set.
@@ 1 Introduction
Differentially private (DP; Dwork et al. (2006)) model training is an effective strategy for privacy-preserving ML on sensitive data. For most optimization-based learning algorithms, DP-SGD (Song et al., 2013; Abadi et al., 2016) can be readily applied to obtain models with rigorous DP guarantee. Regrettably, DP training has also been marred by a sub-optimal privacy-utility trade-off, with model utility severely lagging behind their non-private counterpart (Jayaraman & Evans, 2019; Tramer & Boneh, 2020; Kurakin et al., 2022). At the core of this unfavorable trade-off is the difficulty of DP representation learning. Tramer & Boneh (2020) showed that when DP training from scratch under a low-to-moderate privacy budget, most models do not learn useful representations, with the quality of learned representations worse than even handcrafted features. These observations naturally lead to the research question: “How does one learn useful representations with DP training?”
One plausible reason for the failure of prior attempts at DP representation learning is the lack of training data. Indeed, DP limits the information content of each training sample via the privacy budget $\varepsilon$, inducing a privacy-accuracy-sample size tradeoff; thus a substantially larger training dataset is required to extract the same amount of information to train the model. As the vast majority of prior work only utilize small to moderate scale classification datasets such as CIFAR-10 (Krizhevsky et al., 2009) and ImageNet (Deng et al., 2009b), the amount of training data is simply insufficient for learning high-quality representations under DP (Tramer & Boneh, 2020). Yu et al. (2023) made partial progress towards this through self-supervised learning (SSL) on internet-scale data. By training a masked autoencoder (MAE; He et al. (2022)) using DP-SGD on a 233M subset of the LAION-2B dataset (Schuhmann et al., 2022), the model learned image representations that are on-par with non-private AlexNet (Krizhevsky et al., 2012) trained on ImageNet—the first deep learning model to outperform handcrafted features and a major cornerstone for representation learning. However, the MAE objective also promotes the model to learn extraneous details in the image that may not be helpful for obtaining generalizable representations, severely limiting the potential of this approach for DP.
@@
The image representations learned by DP-Cap also exhibit strong performance for multimodal tasks that require alignment of image and text features, the first occurrence for models trained from scratch with DP; see Figure 1(b). As a qualitative evaluation, we also use the trained DP-Cap model to caption several images from the MS-COCO 2017 (Lin et al., 2015) test set in Figure 1(c) and Appendix B.3. The resulting captions are grammatically correct and semantically coherent, while (close to) accurately describing contents of the image; this is interesting because our model has only been exposed to language supervision from LAION, which are far from being flawless. Our results suggest that DP training on internet-scale multimodal datasets can be a viable approach for obtaining high-utility learned representations.
@@ 2 Background and Related Work
Vision-language pre-training. Many modern ML datasets such as Conceptual Captions (Changpinyo et al., 2021), LAION (Schuhmann et al., 2021) and DataComp (Gadre et al., 2023) consist of aligned image-text pairs where the image and text contain roughly similar semantic information. One can leverage the aligned nature of the training data to pre-train vision-language models (VLMs) that connect the two modalities, whose representations perform more general multi-modal tasks. Contrastive learning-based techniques such as CLIP (Radford et al., 2021) and BLIP (Li et al., 2022a) are also applicable for pre-training VLMs. Doing so not only learns high-quality image and text representations but also introduces new multi-modal capabilities such as cross-modal retrieval and zero-shot prediction (Radford et al., 2021). Recent work by (Tschannen et al., 2023) shows an image captioning approach (predicting text captions from images) is a viable alternative to contrastive learning and can lead to models with robust performance.
@@
${\small\widetilde{\mathbf{g}}_{k}:=\frac{1}{B}\left[\sum_{i\in\mathcal{B}_{k}}\text{clip}_{C}\left(\nabla_{\bm{\theta}}\ell_{i}(\bm{\theta}_{k})\right)+\mathcal{N}\left(0,C^{2}\sigma^{2}\mathbf{I}\right)\right],}$ (2)
where $\ell_{i}(\bm{\theta})$ is the per-sample loss function evaluated at sample $\mathbf{x}_{i}$. We also use the term “DP-SGD” loosely to refer to the category of gradient-based optimization algorithms that operate on the noisy gradient, e.g., Adam (Kingma & Ba, 2014). The privacy analysis of DP-SGD relies on composition of multiple steps. One particularly powerful analysis framework amenable to such compositions relies on a variant of DP called Rényi differential privacy (RDP) (Mironov, 2017). An advantage of RDP is its additive composition property, where the privacy guarantees of a sequence of mechanisms can be combined with amplification by subsampling (Wang et al., 2019) and then translated to $(\varepsilon,\delta)$-DP (Balle et al., 2020; Gopi et al., 2021). In this work, we adopt this accounting technique.
Scaling up DP-SGD training. DP training is a theoretically and empirically proven remedy against unintended training data memorization. Even models with large $\varepsilon$ (e.g., $\varepsilon=100$) can empirically defend against privacy attacks (Carlini et al., 2021; Guo et al., 2023). Despite its great appeal, DP training also carries a significant drawback of large drop in model utility (Abadi et al., 2016; Tramer & Boneh, 2020). For example, the SOTA performance on ImageNet when training from scratch with a DP guarantee of $\varepsilon=8$ is 39.2% (Sander et al., 2023); in comparison, the non-private performance on ImageNet when training from scratch can reach 88% (Touvron et al., 2022) or higher. This degradation in model utility also translates to poorly learned representations, as Tramer & Boneh (2020) showed that even handcrafted features can rival ones learned through DP training.
@@ 3.1 DP Representation Learning via Image Captioning
Why is vision-language pre-training suitable? Given image-text aligned datasets, prior works (Radford et al., 2021; Li et al., 2022a; Tschannen et al., 2023) showed that pre-training using language supervision is an appealing option for non-private representation learning. We hypothesize that this is true for DP representation learning as well. Compared to image-only supervision, language supervision contains a more condensed summary of the image content, allowing the model to ignore irrelevant details such as background and focus on objects of interest and their relationships. This is especially helpful for DP since the model needs to extract as much useful information as possible from each sample given the privacy budget $\varepsilon$. Captioning could thus enhance the privacy-utility-sample size trade-off in DP, considering it requires less information per sample.
In addition, we show that vision-language pre-training supports a very large batch size, much larger than what is typically used in image-only pre-training (Radford et al., 2021; Li et al., 2022a; Yu et al., 2022). This subtle aspect is in fact crucial for reducing the effective noise in DP-SGD (Li et al., 2021), which allows the model parameters to converge to a stable solution with lower training loss (see Section 3.2).
@@
Unlike contrastive learning, the image captioning approach (Sariyildiz et al., 2020; Desai & Johnson, 2021; Tschannen et al., 2023) aligns well with DP-SGD training. Specifically, an image captioner is trained to predict captions based on their corresponding images. The training objective of the image captioner for one image-text pair $[\mathbf{x}^{\text{img}},\mathbf{z}^{\text{text}}]$ is to minimize over $\bm{\theta}:=\{\bm{\theta}_{\textsf{enc}},\bm{\theta}_{\textsf{dec}}\}$ the following loss:
$L_{\text{Cap}}(\bm{\theta}):=\frac{1}{T}\sum_{t=0}^{T-1}\ell_{\text{CE}}\Big{(}{z}^{\text{text}}_{t+1},\varphi\big{(}\mathbf{z}^{\mathrm{img}},\mathbf{z}_{1:t}^{\mathrm{text}};\bm{\theta}_{\textsf{dec}}\big{)}\Big{)},$ (3)
$\mathrm{where}\,\mathbf{z}^{\mathrm{img}}=\underbrace{\psi(\mathbf{x}^{\text{img}};\bm{\theta}_{\textsf{enc}})}_{\text{image embedding}},\,\,\mathbf{z}^{\mathrm{text}}_{1:t}=\underbrace{{z}^{\text{text}}_{1},\ldots,{z}^{\text{text}}_{t}}_{\text{first $t$ tokens}}.\vspace{-0.2in}$
We use $\mathbf{z}^{\text{text}}$ denotes the caption token sequence $\{{z}^{\text{text}}_{1},\ldots,{z}^{\text{text}}_{T}\}$, and we let $\bm{z}_{1:0}^{\mathrm{text}}=\emptyset$. The image captioner consists of two parts: the image encoder $\psi(\cdot;\bm{\theta}_{\textsf{enc}})$ and the text decoder $\varphi(\cdot;\bm{\theta}_{\textsf{dec}})$. The rationale behind the design of equation 3 is that the image encoder maps the input image $\mathbf{x}^{\text{img}}$ to an embedding vector, and the text decoder takes the image embedding $\psi(\mathbf{x}^{\text{img}};\bm{\theta}_{\textsf{enc}})$ and the first $t$ caption tokens $\{{z}^{\text{text}}_{1},\ldots,{z}^{\text{text}}_{t}\}$ as inputs and predicts the next caption token ${z}_{t+1}^{\text{text}}$. Both the encoder and decoder are trained to maximize the log-likelihood of the correct next token. Equation 3 corresponds to the loss function for an image-text pair $[\mathbf{x}^{\text{img}},\mathbf{z}^{\text{text}}]$; summing over all the samples in a batch gives the complete empirical loss in an additive form, which is directly compatible with DP-SGD.
Interpretation of DP bound for image-text pairs. $\varepsilon$-DP bounds the amount of information extracted from each training sample by $\varepsilon$. For DP-Cap, each sample is made of {image + caption}, while ViP (Yu et al., 2023) utilizes {image} only. Consequently, DP-Cap inherently offers an equivalent or better privacy guarantee for each image. To better understand this, we can consider membership inference attacks (Shokri et al., 2017) as an example. Suppose $\varepsilon$-DP upper bounds the success rate of the membership inference attack (when given the image-text pair) against DP-Cap as less or equal to $p$. Then the MIA success rate when given only the image can be at most $p$ since the attacker has strictly less information. This is exactly the upper bound for the success rate of a membership inference attack against ViP.
@@ 3.2 Strategy for Effective DP Training
Figure 2: Impact of synthetic initialization on DP-Cap. The learned image representation benefits substantially from initializing on the Shaders21K dataset. The gap between DP-Cap (random init) and DP-Cap (syn init) can be as large as 24% when evaluated using linear probing on ImageNet.
@@
(b) Performance improves consistently with larger batch size
@@
Figure 3: (a) We fix the effective noise $\sigma/B=5.6\times 10^{-7}$ (corresponding to our (B, $\sigma$) = (1.3M, 0.728)) and show that the loss is remarkably consistent across different batch sizes, allowing us to effectively scale up batch size to improve the SNR. (b) Performance from 4 sets of parameters that provide $\varepsilon=8$, with constant number of steps 5708. From batch size 98k (used in ViP (Yu et al., 2023)), to our 1.3M batch size. In contrast to ViP, DP-Cap successfully leverages the better SNR and learns features that achieve substantially better 10-shot accuracy on ImageNet even compared to a non-private MAE (He et al., 2022) trained on the same dataset (see Appendix A.1).
@@
Sufficient number of training steps. We address this challenge via synthetic pre-training. Image representations learned by DP-Cap (random init) outperform those of ViP (random init) as evidenced in table 9 in Appendix B. Interestingly,  Yu et al. (2023) and  Tang et al. (2024) have shown that synthetic images consisting of only textures can provide a good initialization for training DP models without any privacy risk. With this initialization in  Yu et al. (2023), the reconstruction-based model can then focus on learning dataset-specific properties rather than low-level image properties such as edge detectors, therefore expending privacy budget in a more optimal manner. We adapt this technique of pre-training on the Shaders21K dataset (Baradad et al., 2022) for initialization and observe that it is even more effective with DP-Cap compared to ViP. [^1] For the synthetic pre-training on Shders21k in DP-Cap, we modify the training of image captioning model such that the model is trained to predict missing image patches, which is similar to MAE (He et al., 2022). More details can be found in Appendix A.1. As shown in Fig. 2, our DP-Cap (syn init) improves over DP-Cap (random init) by more than 24% on ImageNet-1k linear probing. It also improves over synthetic initialization alone (Syn-init) by more than 14% on ImageNet-1k linear probing, whereas the gain in ViP is smaller than 6%.
@@
Using extreme batch sizes to reduce effective noise. Li et al. (2021) first showed that increasing the batch size in DP-SGD often improves the privacy-utility trade-off. This is because the effective noise added to the average gradient has magnitude $\sigma/B$ (cf. equation 2), and that increasing $B$, rather than decreasing $\sigma$, results in better privacy guarantees according to conventional accounting techniques (Bun & Steinke, 2016; Dwork & Rothblum, 2016; Mironov et al., 2019; Sander et al., 2023). However, in supervised learning, increasing the batch size beyond a certain point may degrade performance in both private and non-private settings (Sander et al., 2024). Specifically, Sander et al. (2023) observed that when training a classifier from scratch with DP-SGD on ImageNet, at a fixed number of steps $S$ and fixed effective noise $\sigma/B$, the performance decreases significantly when the batch size becomes too large: for $B\in[128,16384]$, a drop of 10% in top-1 accuracy was observed.
@@
Intriguingly, we find that vision-language pre-training on internet-scale datasets can tolerate extreme batch sizes, e.g. $B=1$M. In Figure 3(a), we compare the loss behaviors when scaling the batch size for DP-Cap. We fix the effective noise $\sigma/B$ while varying the batch size. In stark contrast to the previous observation from (Sander et al., 2023), the loss trajectory is identical across different batch sizes. With this observation, we are able to successfully scale up the batch size for DP-Cap to as large as $B=1.3$M, achieving an effective noise of $5.6\times 10^{-7}$, almost 10 times smaller than the effective noise of ViP in Yu et al. (2023). Training DP-Cap under such a small effective noise allows it to extract information from the training dataset more efficiently under the DP constraint. In Figure 3, we show that with $B=1.3$M, the representation learned by DP-Cap even outperforms non-private MAE trained on the same dataset. From these results, we find that the DP-Cap training can handle extreme large batches much more effectively than ViP in the context of differentially private training. This advantage is likely independent of image captioning models’s inherent superiority over MAE.
@@
Figure 4: Number of GPU hours to train DP-Cap for a single epoch on 233M samples. For the Large model, we achieve a close to 5×5$\times$5 × reduction.
Method Model
Tiny Small Base Large
functorch 36 20 12 5
Ghost norm 350 230 180 64
Ghost norm+AMP 420 310 225 90
@@
Table 1: Max physical batch size under different per-sample gradient computation methods.
@@
Improving training pipeline efficiency. DP training is known to be computationally inefficient due to factors such as per-sample gradient computation (Lee & Kifer, 2020; Li et al., 2021). Training on internet-scale datasets using extreme batch sizes further complicates this issue. For example, a naive implementation of DP-Cap with per-sample gradient computation using functorch would take approximately 61 days(!) on 128 NVIDIA V100 GPUs. Recent studies (Li et al., 2021; Bu et al., 2023) have enhanced this efficacy by enabling DP training without the necessity to store per-sample gradients. We made significant efficiency improvements to the training pipeline using two techniques: ghost norm (Li et al., 2021) and the automatic mixed precision (AMP) package in PyTorch. Combining these two techniques with DP-SGD requires careful considerations to ensure both a correct DP guarantee as well as numerical stability. We detail the implementation in Appendix A.2.